OpenStudy (anonymous):

HELPPP!! Find the dimensions of a box with a square base with (a) volume 12 and the minimal surface area (b) surface area 20 and minimal volume

OpenStudy (amistre64):

The rate of change of volume is surface area ... in other words, as fast as the surface area changes, so does the volume

OpenStudy (amistre64):

Sa = 4(b*h) + 2(b*b) assuming its a solid box with 4 sides and a top and bottom

OpenStudy (amistre64):

V = b*b*h 12 = b^2 h h = 12/b^2

OpenStudy (amistre64):

Sa = 4(b*h) + 2(b*b) Sa = 4(b*12/b^2) + 2b^2 Sa = 48/b + 2b^2 we can find min or max with the first derivative Sa' = -48/b^2 + 4b Sa' = -48 + 4b^3/b^2 = 0 when: -48 + 4b^3 = 0 -12 + b^3 = 0 b = cbrt(12) , maybe